# Loss of integrability in a system with two-photon interactions

### v1

@@ III.3 Breaking the integrability
We investigate the correspondence between the integrable case ($\omega_{0}=0$), described by Eq. (6), and the non-integrable case ($\omega_{0}\neq 0$), represented by Eq. (1). As a first step, we focus on the effects of small atomic level splittings on the system’s integrability and examine how deviations from the integrable limit influence the behavior of the observables.
@@
In order to explore the observed regularities, in Fig. 3, we plot the Peres lattice of the operators $\hat{a}^{\dagger}\hat{a}$, $\hat{J}_{z}$, and $\hat{J}_{x}^{2}$. In Fig. 3(a), we highlight the expectation values that appear horizontally compressed in the Peres lattice of $\hat{a}^{\dagger}\hat{a}$, marking them in green. The remaining points, which are arranged along parallel diagonal lines, are shown in blue. This color scheme is then consistently applied to Fig. 3(b) and Fig. 3(c). The green points, originally associated with horizontally contracted structures in the Peres lattice of $\hat{a}^{\dagger}\hat{a}$, exhibit a dependence with the energy which is close to linear for $\hat{J}_{z}$, and parabolic for $\hat{J}_{x}^{2}$. On the other hand, the blue points, which were diagonally grouped in Fig. 3(a), show a linear dependence with the energy for $\hat{J}_{x}^{2}$, and a curved one for $\hat{J}_{z}$.
The last behaviors can be easily understood for the lowest energy eigenstates. As shown in Fig. 3(a), the horizontal set of points, colored on green, have zero photons, and are very close to be eigenstates of $\hat{J}_{z}$, with energies $\omega_{0}m_{z},-j\leq m_{z}\leq j$, as can be seen in the lowest energy region of Fig. 3(b). For these states, $\langle j,m_{z}|\hat{J}_{x}^{2}|j,m_{z}\rangle=(j(j+1)-m_{z}^{2})/2$. The last expression describes the parabolic form observed in the first set of green points in Fig. 3(c), whose values run from $\langle\hat{J}_{x}^{2}\rangle/j^{2}=1/(2j)$ for $m_{z}=\pm j$ to $\langle\hat{J}_{x}^{2}\rangle/j^{2}=(1+1/j)/2$ for $m_{z}=0$.
@@ IV.1 Classical and quantum chaos
Figure 4: [(a1)-(e1)] Average spectral ratio $\langle r\rangle$ and Peres lattices of the operators [(a2)-(e2)] $\hat{a}^{\dagger}\hat{a}$, [(a3)-(e3)] $\hat{J}_{z}$, and [(a4)-(e4)] $\hat{J}_{x}^{2}$ covering a wide energy spectrum. Each column identifies a different coupling strength: [(a1)-(a4)] $\gamma=0.01$, [(b1)-(b4)] $\gamma=0.05$, [(c1)-(c4)] $\gamma=0.1$, [(d1)-(d4)] $\gamma=0.2$, and [(e1)-(e4)] $\gamma=0.3$. In panels (a1)-(e1), the black solid line represents the average spectral ratio in Eq. (16) using the converged eigenvalues of the two-photon Dicke Hamiltonian in Eq. (1). We average the spectral ratio over the four parity subspaces to smooth spectral fluctuations. For each panel from (a1) to (e1), we get around $2\times 10^{4}$ converged eigenvalues for each subspace. The red (blue) dashed line represents the analytical average spectral ratio of the GOE (Poisson) distribution. In panels (a2)-(e2), (a3)-(e3), and (a4)-(e4), the blue dots represent the numerical expectation values using the eigenstates of the two-photon Dicke Hamiltonian for the parity subspace with eigenvalue $p=1$. System parameters: $\omega=1$, $\omega_{0}=\pi/5$, and $j=25$. The truncation value is $n_{\text{max}}=2\times 10^{3}$.
@@
Figure 5: [(a1)-(e1)] Average spectral ratio $\langle r\rangle$ and Peres lattices of the operators [(a2)-(e2)] $\hat{a}^{\dagger}\hat{a}$, [(a3)-(e3)] $\hat{J}_{z}$, and [(a4)-(e4)] $\hat{J}_{x}^{2}$ covering a wide energy spectrum. Each column identifies a different coupling strength: [(a1)-(a4)] $\gamma=0.01$, [(b1)-(b4)] $\gamma=0.05$, [(c1)-(c4)] $\gamma=0.1$, [(d1)-(d4)] $\gamma=0.2$, and [(e1)-(e4)] $\gamma=0.3$. In panels (a1)-(e1), the black solid line represents the average spectral ratio in Eq. (16) using the converged eigenvalues of the two-photon Dicke Hamiltonian in Eq. (1). We average the spectral ratio over the four parity subspaces to smooth spectral fluctuations. For each panel from (a1) to (e1), we get around $2\times 10^{4}$ converged eigenvalues for each subspace. The red (blue) dashed line represents the analytical average spectral ratio of the GOE (Poisson) distribution. In panels (a2)-(e2), (a3)-(e3), and (a4)-(e4), the green dots represent the numerical expectation values using the eigenstates of the two-photon Dicke Hamiltonian for the parity subspace with eigenvalue $p=1$. System parameters: $\omega=1$, $\omega_{0}=2$, and $j=25$. The truncation value is $n_{\text{max}}=2\times 10^{3}$.
@@
Figure 6: [(a1)-(e1)] Average spectral ratio $\langle r\rangle$ and Peres lattices of the operators [(a2)-(e2)] $\hat{a}^{\dagger}\hat{a}$, [(a3)-(e3)] $\hat{J}_{z}$, and [(a4)-(e4)] $\hat{J}_{x}^{2}$ covering a wide energy spectrum. Each column identifies a different atomic level splitting: [(a1)-(a4)] $\omega_{0}=0.01$, [(b1)-(b4)] $\omega_{0}=0.2$, [(c1)-(c4)] $\omega_{0}=0.5$, [(d1)-(d4)] $\omega_{0}=1$, and [(e1)-(e4)] $\omega_{0}=2$. In panels (a1)-(e1), the black solid line represents the average spectral ratio in Eq. (16) using the converged eigenvalues of the two-photon Dicke Hamiltonian in Eq. (1). We average the spectral ratio over the four parity subspaces to smooth spectral fluctuations. For each panel from (a1) to (e1), we get around $2\times 10^{4}$ converged eigenvalues for each subspace. The red (blue) dashed line represents the analytical average spectral ratio of the GOE (Poisson) distribution. In panels (a2)-(e2), (a3)-(e3), and (a4)-(e4), the pink dots represent the numerical expectation values using the eigenstates of the two-photon Dicke Hamiltonian for the parity subspace with eigenvalue $p=1$. In column [(e1)-(e4)], the vertical dashed lines represent the classical energy shells used in Fig. 7. System parameters: $\omega=1$, $\gamma=0.3$, and $j=25$. The truncation value is $n_{\text{max}}=2\times 10^{3}$.
@@
Figure 7: [(a1)-(e1)] Poincaré sections projected on the atomic phase space of the classical two-photon Dicke Hamiltonian in Eq. (5). Each panel identifies a different classical energy shell: (a1) $\epsilon=-1$, (b1) $\epsilon=0$, (c1) $\epsilon=1$, (d1) $\epsilon=5$, and (e1) $\epsilon=10$. The black solid circle represents the available phase space of the selected energy shell. We choose initial conditions distributed over the phase space with $q_{0}=q_{+}(\epsilon,p_{0}=0,Q_{0},P_{0})$. The selected classical energy shells are shown in column [(e1)-(e4)] of Fig. 6 as vertical dashed lines. [(a2)-(e2)] Nearest-neighbor spacing distributions $P(s)$ for the converged eigenvalues of the two-photon Dicke Hamiltonian in Eq. (1). In each panel, we use the converged eigenvalues contained in an energy window with mean $\langle\epsilon\rangle=\langle E\rangle/j$ and width $\sigma_{\epsilon}$: (a2) $\langle\epsilon\rangle=-1.01$ and $\sigma_{\epsilon}=0.10$, (b2) $\langle\epsilon\rangle=-0.08$ and $\sigma_{\epsilon}=0.55$, (c2) $\langle\epsilon\rangle=0.98$ and $\sigma_{\epsilon}=0.34$, (d2) $\langle\epsilon\rangle=5.00$ and $\sigma_{\epsilon}=0.26$, and (e2) $\langle\epsilon\rangle=10.00$ and $\sigma_{\epsilon}=1.35$. We use the parity subspace with eigenvalue $p=1$ and use different system sizes to get an appropriate number of eigenvalues inside the energy window: (a2) $j=100$ and 461 eigenvalues, [(b2)-(d2)] $j=50$ and 1201 eigenvalues, and (e2) $j=25$ and 1601 eigenvalues. The red solid (blue dashed) line represents the analytical GOE (Poisson) distribution in Eq. (14) [Eq. (15)]. Each panel shows the Anderson-Darling parameter who ensures the agreement between the numerical data and the analytical distributions. System parameters: $\omega=1$, $\omega_{0}=2$, and $\gamma=0.3$. The truncation value is (a2) $n_{\text{max}}=500$, [(b2)-(d2)] $n_{\text{max}}=900$, and (e2) $n_{\text{max}}=2\times 10^{3}$.
@@ IV.2 Integrability and coupling strength
In this section, we investigate the impact of the coupling strength $\gamma$ on the integrability of the system, focusing on values below the critical threshold at which spectral collapse occurs ($\gamma<\omega/2$). Our analysis is confined to the normal phase, characterized by a vanishing average photon number. We explore the breakdown of integrability across the energy spectrum, from the ground state to high-energy regions. To quantify the degree of spectral correlations, we employ the spectral ratio defined in Eq. (16), averaging it over finite energy windows to distinguish between GOE and Poisson statistics, thereby identifying signatures of quantum chaos or regular behavior in the system. We analyze two representative cases, one with a small atomic level splitting and one with a large value.
IV.2.1 Small atomic level splitting: $\omega_{0}=\pi/5$
@@
In Figs. 4(a1)–4(e1), we present the average spectral ratio $\langle r\rangle$ as a function of increasing coupling strength $\gamma$. To obtain these results, we diagonalize the two-photon Dicke Hamiltonian following the procedure outlined in Appendix C, using a small atomic level splitting of $\omega_{0}=\pi/5\approx 0.628$. The energy spectra from all four parity subspaces are computed, and the spectral ratio is calculated within each subspace by averaging over finite energy windows. To mitigate strong spectral fluctuations and enhance the robustness of the analysis, we subsequently average the $\langle r\rangle$ values in the four subspaces.
@@
The remaining rows in Fig. 4 display the Peres lattices for various observables corresponding to the same set of increasing coupling strengths. Figures 4(a2)–4(e2) show the Peres lattices of the photon number operator $\hat{a}^{\dagger}\hat{a}$, while Figs. 4(a3)–4(e3) and Figs. 4(a4)–4(e4) display those of the pseudospin operators $\hat{J}_{z}$ and $\hat{J}_{x}^{2}$, respectively. For clarity in visualizing the structure of the Peres lattices, we restrict this part of the analysis to a single-parity subspace with eigenvalue $p=1$.
@@
Overall, the coupling strength $\gamma$ plays a key role in driving a transition from the integrable to the non-integrable regime of the system. This transition is clearly reflected in the increasing disorder of the Peres lattices as $\gamma$ increases. The loss of integrability is particularly evident in the Peres lattices of the operator $\hat{J}_{x}^{2}$ [see Figs. 4(a4)–4(e4)], where the initially ordered structures progressively give way to irregular, chaotic patterns. Notably, even in those lattices that exhibit significant disorder at higher energies, ordered patterns persist in the low-energy region [see Figs. 4(d2)–4(d4) and Figs. 4(e2)–4(e4), column view], indicating that the system retains integrable characteristics near the ground state.
When analyzing the behavior of the average spectral ratio $\langle r\rangle$, we observe in Fig. 4(a1) that, for the smallest coupling strength $\gamma=0.01$, there are significant deviations from the expected Poisson statistics. Despite the corresponding Peres lattices displaying well-ordered structures [see Figs. 4(a2)–4(a4), column view], the spectral ratio fails to accurately reflect the integrable nature of the system. The same behavior is observed in Fig. 4(b1), $\langle r\rangle$ initially takes on large values in the low-energy region and then decays toward values consistent with Poisson statistics at intermediate energies. These elevated values of $\langle r\rangle$ are characteristic of nearly harmonic spectra, such as that of the harmonic oscillator, for which $\langle r\rangle=1$. This observation is consistent with the nearly equidistant lines representing the expectation values of $\hat{J}_{z}$ in Fig. 4(a3), and in the low-energy region of Fig. 4(b3). In Fig. 4(b4), at energies $\epsilon>35$ the energy eigenstates tend to be also eigenstates of $\hat{J}_{x}$, opening along the different values of $m_{x}^{2}$, which also explains that the expectation values of $\hat{J}_{z}$ in Fig. 4(b3) tend to zero.
A similar initial trend in $\langle r\rangle$ is observed in Fig. 4(c1), where it has large values at low energies before converging toward the Poisson limit. However, at higher energies, $\langle r\rangle$ begins to rise again, signaling the onset of spectral correlations. This behavior aligns with the increasing disorder observed in the corresponding Peres lattices [see Figs. 4(c2)–4(c4), column view], marking the beginning of a transition toward non-integrability.
Figure 4(d1) and Fig. 4(e1) show that the average spectral ratio $\langle r\rangle$ fluctuates around the value expected for GOE statistics in certain energy regions. This behavior is consistent with the strong disorder observed in the corresponding Peres lattices at high energies [see Figs. 4(d2)–4(d4) and Figs. 4(e2)–4(e4), column view]. As the coupling strength $\gamma$ increases, a clear trend emerges in which the spectral ratio transitions from values characteristic of Poisson statistics to those indicative of GOE statistics, reflecting a progressive loss of integrability and the emergence of quantum chaos.
With the exception of the case $\gamma=0.01$, where the energy spectrum remains nearly harmonic and thus obscures the spectral statistics, we observe a strong correspondence between the degree of disorder in the Peres lattices and the values of the average spectral ratio $\langle r\rangle$. This agreement reinforces the complementary nature of these quantum chaos indicators and underscores the spectral ratio’s usefulness as a diagnostic tool for integrability breakdown in the two-photon Dicke model.
IV.2.2 Large atomic level splitting: $\omega_{0}=2$
We now extend the previous analysis to a larger value of the atomic level splitting, $\omega_{0}=2$. In Fig. 5, we present the same set of diagnostics as before, the average spectral ratio $\langle r\rangle$ and the Peres lattices corresponding to the operators $\hat{a}^{\dagger}\hat{a}$, $\hat{J}_{z}$, and $\hat{J}_{x}^{2}$. These are shown in Figs. 5(a1)–5(e1), Figs. 5(a2)–5(e2), Figs. 5(a3)–5(e3), and Figs. 5(a4)–5(e4), respectively.
As in the previous case with $\omega_{0}=\pi/5$, we observe for the case with $\omega_{0}=2$ a consistent trend in which increasing the coupling strength $\gamma$ drives a transition from integrable to non-integrable behavior. This transition is again clearly reflected in the increasing disorder of the Peres lattices, particularly for the operator $\hat{J}_{x}^{2}$ [see Figs. 5(a4)–5(e4)]. The persistence of integrable dynamics near the ground state is newly detected for the low-energy region of the Peres lattices that exhibit strong disorder at higher energies [see Figs. 5(d2)–5(d4) and Figs. 5(e2)–5(e4), column view].
@@ IV.2.2 Large atomic level splitting: $\omega_{0}=2$
The behavior of the average spectral ratio $\langle r\rangle$ with $\omega_{0}=2$ also exhibits a trend similar to that observed in the previous case with $\omega_{0}=\pi/5$. In Figs. 5(a1)–5(c1), $\langle r\rangle$ shows a decay from large values at low energies, followed by fluctuations around the value associated with Poisson statistics. Notably, in Fig. 5(c1), the spectral ratio increases again at higher energies, approaching values characteristic of GOE statistics. This shift is consistent with the emergence of disorder in the corresponding Peres lattices at high energies [see Figs. 5(c2)–5(c4), column view], further indicating the onset of non-integrable dynamics in that energy regime. At these relatively low values of $\gamma$, we observe peculiar sets of regular lines in the expectation values of $\hat{J}_{x}^{2}$ at high energies, $\epsilon\approx 40$ in Fig. 5(b4) and $\epsilon\approx 20$ in Fig. 5(c4). These unexpectedly ordered structures in the middle of an otherwise disordered regime suggest underlying mechanisms that warrant further investigation.
@@
The final panels, Fig. 5(d1) and Fig. 5(e1), show that the average spectral ratio $\langle r\rangle$ stabilizes around the GOE limit at high energies, following a transition from Poisson-like statistics at lower energies. This behavior aligns with the corresponding Peres lattices, which display pronounced disorder in the high-energy regime and ordered structures at low energies [see Figs. 5(d2)–5(d4) and Figs. 5(e2)–5(e4), column view]. Overall, for the present case with $\omega_{0}=2$, we again observe a clear transition from Poisson to GOE statistics and from ordered to disordered Peres lattices as the coupling strength $\gamma$ increases from low to high values.

### v2

@@ III.3 Breaking the integrability
We investigate the correspondence between the integrable case ($\omega_{0}=0$), described by Eq. (7), and the nonintegrable case ($\omega_{0}\neq 0$), represented by Eq. (1). As a first step, we focus on the effects of small atomic level splittings on the system’s integrability and examine how deviations from the integrable limit influence the behavior of the observables.
@@
To explore the observed regularities, in Fig. 3, we plot the Peres lattice of the operators $\hat{a}^{\dagger}\hat{a}$, $\hat{J}_{z}$, and $\hat{J}_{x}^{2}$. In Fig. 3(a), we highlight the expectation values that appear horizontally compressed in the Peres lattice of $\hat{a}^{\dagger}\hat{a}$, marking them in green. The remaining points, which are arranged along parallel diagonal lines, are shown in blue. This color scheme is then consistently applied to Fig. 3(b) and Fig. 3(c). The green points, originally associated with horizontally contracted structures in the Peres lattice of $\hat{a}^{\dagger}\hat{a}$, exhibit a dependence with the energy which is close to linear for $\hat{J}_{z}$, and parabolic for $\hat{J}_{x}^{2}$. On the other hand, the blue points, which were diagonally grouped in Fig. 3(a), show a linear dependence with the energy for $\hat{J}_{x}^{2}$, and a curved one for $\hat{J}_{z}$.
The last behaviors can be easily understood for the lowest-energy eigenstates. As shown in Fig. 3(a), the horizontal set of points, colored on green, have zero photons, and are very close to be eigenstates of $\hat{J}_{z}$, with energies $\omega_{0}m_{z},-j\leq m_{z}\leq j$, as can be seen in the lowest-energy region of Fig. 3(b). For these states, $\langle j,m_{z}|\hat{J}_{x}^{2}|j,m_{z}\rangle=[j(j+1)-m_{z}^{2}]/2$. The last expression describes the parabolic form observed in the first set of green points in Fig. 3(c), whose values run from $\langle\hat{J}_{x}^{2}\rangle/j^{2}=1/(2j)$ for $m_{z}=\pm j$ to $\langle\hat{J}_{x}^{2}\rangle/j^{2}=(1+1/j)/2$ for $m_{z}=0$.
@@ IV.1 Classical and quantum chaos
Figure 4: (a1)-(e1) Average spectral ratio $\langle r\rangle$ and Peres lattices of the operators (a2)-(e2) $\hat{a}^{\dagger}\hat{a}$, (a3)-(e3) $\hat{J}_{z}$, and (a4)-(e4) $\hat{J}_{x}^{2}$ covering a wide energy spectrum, where $\epsilon=E/j$. Each column identifies a different coupling strength: (a1)-(a4) $\gamma=0.01$, (b1)-(b4) $\gamma=0.05$, (c1)-(c4) $\gamma=0.1$, (d1)-(d4) $\gamma=0.2$, and (e1)-(e4) $\gamma=0.3$. In panels (a1)-(e1), the black solid line represents the average spectral ratio in Eq. (17) using the converged eigenvalues of the two-photon Dicke Hamiltonian in Eq. (1). We average the spectral ratio over the four parity subspaces to smooth spectral fluctuations. For each panel from (a1) to (e1), we get around $2\times 10^{4}$ converged eigenvalues for each subspace. The red (blue) dashed line represents the analytical average spectral ratio of the GOE (Poisson) distribution. In panels (a2)-(e2), (a3)-(e3), and (a4)-(e4), the green dots represent the numerical expectation values using the eigenstates of the two-photon Dicke Hamiltonian for the parity subspace with eigenvalue $p=1$. System parameters: $\omega=1$, $\omega_{0}=2$, and $j=25$. The truncation value is $n_{\text{max}}=2\times 10^{3}$.
@@
Figure 5: (a1)-(e1) Average spectral ratio $\langle r\rangle$ and Peres lattices of the operators (a2)-(e2) $\hat{a}^{\dagger}\hat{a}$, (a3)-(e3) $\hat{J}_{z}$, and (a4)-(e4) $\hat{J}_{x}^{2}$ covering a wide energy spectrum, where $\epsilon=E/j$. Each column identifies a different coupling strength: (a1)-(a4) $\gamma=0.01$, (b1)-(b4) $\gamma=0.05$, (c1)-(c4) $\gamma=0.1$, (d1)-(d4) $\gamma=0.2$, and (e1)-(e4) $\gamma=0.3$. In panels (a1)-(e1), the black solid line represents the average spectral ratio in Eq. (17) using the converged eigenvalues of the two-photon Dicke Hamiltonian in Eq. (1). We average the spectral ratio over the four parity subspaces to smooth spectral fluctuations. For each panel from (a1) to (e1), we get around $2\times 10^{4}$ converged eigenvalues for each subspace. The red (blue) dashed line represents the analytical average spectral ratio of the GOE (Poisson) distribution. In panels (a2)-(e2), (a3)-(e3), and (a4)-(e4), the blue dots represent the numerical expectation values using the eigenstates of the two-photon Dicke Hamiltonian for the parity subspace with eigenvalue $p=1$. System parameters: $\omega=1$, $\omega_{0}=\pi/5$, and $j=25$. The truncation value is $n_{\text{max}}=2\times 10^{3}$.
@@
Figure 6: (a1)-(e1) Average spectral ratio $\langle r\rangle$ and Peres lattices of the operators (a2)-(e2) $\hat{a}^{\dagger}\hat{a}$, (a3)-(e3) $\hat{J}_{z}$, and (a4)-(e4) $\hat{J}_{x}^{2}$ covering a wide energy spectrum, where $\epsilon=E/j$. Each column identifies a different atomic level splitting: (a1)-(a4) $\omega_{0}=0.01$, (b1)-(b4) $\omega_{0}=0.2$, (c1)-(c4) $\omega_{0}=0.5$, (d1)-(d4) $\omega_{0}=1$, and (e1)-(e4) $\omega_{0}=2$. In panels (a1)-(e1), the black solid line represents the average spectral ratio in Eq. (17) using the converged eigenvalues of the two-photon Dicke Hamiltonian in Eq. (1). We average the spectral ratio over the four parity subspaces to smooth spectral fluctuations. For each panel from (a1) to (e1), we get around $2\times 10^{4}$ converged eigenvalues for each subspace. The red (blue) dashed line represents the analytical average spectral ratio of the GOE (Poisson) distribution. In panels (a2)-(e2), (a3)-(e3), and (a4)-(e4), the pink dots represent the numerical expectation values using the eigenstates of the two-photon Dicke Hamiltonian for the parity subspace with eigenvalue $p=1$. In column (e1)-(e4), the vertical dashed lines represent the classical energy shells used in Fig. 7. System parameters: $\omega=1$, $\gamma=0.3$, and $j=25$. The truncation value is $n_{\text{max}}=2\times 10^{3}$.
Figure 7: (a1)-(e1) Poincaré sections projected on the atomic phase space of the classical two-photon Dicke Hamiltonian in Eq. (5). Each panel identifies a different classical energy shell: (a1) $\epsilon=-1$, (b1) $\epsilon=0$, (c1) $\epsilon=1$, (d1) $\epsilon=5$, and (e1) $\epsilon=10$. The black solid circle represents the available phase space of the selected energy shell. We choose initial conditions distributed over the phase space with $q_{0}=q_{+}(\epsilon,p_{0}=0,Q_{0},P_{0})$. The selected classical energy shells are shown in column (e1)-(e4) of Fig. 6 as vertical dashed lines. (a2)-(e2) Nearest-neighbor spacing distributions $P(s)$ for the converged eigenvalues of the two-photon Dicke Hamiltonian in Eq. (1). In each panel, we use the converged eigenvalues contained in an energy window with mean $\langle\epsilon\rangle=\langle E\rangle/j$ and width $\sigma_{\epsilon}$: (a2) $\langle\epsilon\rangle=-1.01$ and $\sigma_{\epsilon}=0.10$, (b2) $\langle\epsilon\rangle=-0.08$ and $\sigma_{\epsilon}=0.55$, (c2) $\langle\epsilon\rangle=0.98$ and $\sigma_{\epsilon}=0.34$, (d2) $\langle\epsilon\rangle=5.00$ and $\sigma_{\epsilon}=0.26$, and (e2) $\langle\epsilon\rangle=10.00$ and $\sigma_{\epsilon}=1.35$. We use the parity subspace with eigenvalue $p=1$ and use different system sizes to get an appropriate number of eigenvalues inside the energy window: (a2) $j=100$ and 461 eigenvalues, (b2)-(d2) $j=50$ and 1201 eigenvalues, and (e2) $j=25$ and 1601 eigenvalues. The red solid (blue dashed) line represents the analytical GOE (Poisson) distribution in Eq. (15) [Eq. (16)]. Each panel shows the Anderson-Darling parameter, which ensures the agreement between the numerical data and the analytical distributions. System parameters: $\omega=1$, $\omega_{0}=2$, and $\gamma=0.3$. The truncation value is (a2) $n_{\text{max}}=500$, (b2)-(d2) $n_{\text{max}}=900$, and (e2) $n_{\text{max}}=2\times 10^{3}$.
@@ IV.2 Integrability and coupling strength
In this section, we investigate the impact of the coupling strength $\gamma$ on the integrability of the system, focusing on values below the critical threshold at which spectral collapse occurs ($\gamma<\omega/2$). Our analysis is confined to the normal phase, characterized by a vanishing average photon number. We explore the breakdown of integrability across the energy spectrum, from the ground state to high-energy regions. To quantify the degree of spectral correlations, we employ the spectral ratio defined in Eq. (17), averaging it over finite-energy windows to distinguish between GOE and Poisson statistics, thereby identifying signatures of quantum chaos or regular behavior in the system. We analyze two representative cases, one where the field frequency and the atomic level splitting are in resonance and one out of resonance.
IV.2.1 Resonant case: $\omega_{0}=2\omega$
@@
In Figs. 4(a1)–4(e1), we present the average spectral ratio $\langle r\rangle$ as a function of increasing coupling strength $\gamma$. To obtain these results, we diagonalize the two-photon Dicke Hamiltonian following the procedure outlined in Appendix C, using an atomic level splitting in resonance with the field frequency, $\omega=1$ and $\omega_{0}=2\omega=2$. The energy spectra from all four parity subspaces are computed, and the spectral ratio is calculated within each subspace by averaging over finite-energy windows. To mitigate strong spectral fluctuations and enhance the robustness of the analysis, we subsequently average the $\langle r\rangle$ values in the four subspaces.
@@
The remaining rows in Fig. 4 display the Peres lattices for various observables corresponding to the same set of increasing coupling strengths. Figures 4(a2)–4(e2) show the Peres lattices of the photon number operator $\hat{a}^{\dagger}\hat{a}$, while Figs. 4(a3)–4(e3) and 4(a4)–4(e4) display those of the pseudospin operators $\hat{J}_{z}$ and $\hat{J}_{x}^{2}$, respectively. For clarity in visualizing the structure of the Peres lattices, we restrict this part of the analysis to a single-parity subspace with eigenvalue $p=1$.
@@
Overall, the coupling strength $\gamma$ plays a key role in driving a transition from the integrable to the nonintegrable regime of the system. This transition is clearly reflected in the increasing disorder of the Peres lattices as $\gamma$ increases. The loss of integrability is particularly evident in the Peres lattices of the operator $\hat{J}_{x}^{2}$ [see Figs. 4(a4)–4(e4)], where the initially ordered structures progressively give way to irregular, chaotic patterns. Notably, even in those lattices that exhibit significant disorder at higher energies, ordered patterns persist in the low-energy region [see Figs. 4(d2)–4(d4) and 4(e2)–4(e4), column view], indicating that the system retains integrable characteristics near the ground state.
@@
In Figs. 4(a1)–4(c1), the average spectral ratio $\langle r\rangle$ shows a decay from large values at low energies, followed by fluctuations around the value associated with Poisson statistics. These elevated values of $\langle r\rangle$ are characteristic of nearly harmonic spectra, such as that of the harmonic oscillator, for which $\langle r\rangle=1$. Notably, in Fig. 4(c1), the spectral ratio increases again at higher energies, approaching values characteristic of GOE statistics. This shift is consistent with the emergence of disorder in the corresponding Peres lattices at high energies [see Figs. 4(c2)–4(c4), column view], further indicating the onset of nonintegrable dynamics in that energy regime. At these relatively low values of $\gamma$, we observe peculiar sets of regular lines in the expectation values of $\hat{J}_{x}^{2}$ at high energies, $\epsilon\approx 40$ in Fig. 4(b4) and $\epsilon\approx 20$ in Fig. 4(c4). These unexpectedly ordered structures in the middle of an otherwise disordered regime suggest underlying mechanisms that warrant further investigation.
@@
The final panels, Fig. 4(d1) and 4(e1) show that the average spectral ratio $\langle r\rangle$ stabilizes around the GOE limit at high energies, following a transition from Poisson-like statistics at lower energies. This behavior aligns with the corresponding Peres lattices, which display pronounced disorder in the high-energy regime and ordered structures at low energies [see Figs. 4(d2)–4(d4) and 4(e2)–4(e4), column view].
In general, for the resonant case with $\omega_{0}=2\omega$, we observe a clear transition from Poisson to GOE statistics and from ordered to disordered Peres lattices as the coupling strength $\gamma$ increases from low to high values. This agreement reinforces the complementary nature of these quantum chaos indicators and underscores the spectral ratio’s usefulness as a diagnostic tool for integrability breakdown in the two-photon Dicke model.
IV.2.2 Nonresonant case: $\omega_{0}=(\pi/5)\omega$
We now extend the previous analysis to an atomic level splitting out of resonance with the field frequency, $\omega=1$ and $\omega_{0}=(\pi/5)\omega\approx 0.628$. In Fig. 5, we present the same set of diagnostics as before, the average spectral ratio $\langle r\rangle$ and the Peres lattices corresponding to the operators $\hat{a}^{\dagger}\hat{a}$, $\hat{J}_{z}$, and $\hat{J}_{x}^{2}$. These are shown in Figs. 5(a1)–5(e1),  5(a2)–5(e2),  5(a3)–5(e3), and 5(a4)–5(e4), respectively.
As in the resonant case with $\omega_{0}=2\omega$, we observe for the nonresonant case with $\omega_{0}=(\pi/5)\omega$ a consistent trend in which increasing the coupling strength $\gamma$ drives a transition from integrable to nonintegrable behavior. This transition is again clearly reflected in the increasing disorder of the Peres lattices, particularly for the operator $\hat{J}_{x}^{2}$ [see Figs. 5(a4)–5(e4)]. The persistence of integrable dynamics near the ground state is newly detected for the low-energy region of the Peres lattices that exhibit strong disorder at higher energies [see Figs. 5(d2)–5(d4) and 5(e2)–5(e4), column view].
The behavior of the average spectral ratio $\langle r\rangle$ for the nonresonant case with $\omega_{0}=(\pi/5)\omega$ also exhibits an overall trend similar to that observed in the resonant case with $\omega_{0}=2\omega$. A special case is first illustrated in Fig. 5(a1) for the smallest coupling strength $\gamma=0.01$, where significant deviations from the expected Poisson statistics are observed. Despite the corresponding Peres lattices displaying well-ordered structures [see Figs. 5(a2)–5(a4), column view], the spectral ratio fails to accurately reflect the integrable nature of the system.
As occurred in the resonant case [see Figs. 4(a1)-4(c1)], a similar behavior is observed for the nonresonant case in Fig. 5(b1), where $\langle r\rangle$ initially takes on large values in the low-energy region and then decays toward values consistent with Poisson statistics at intermediate energies. This observation is consistent with the nearly equidistant lines representing the expectation values of $\hat{J}_{z}$ in Fig. 5(a3), and in the low-energy region of Fig. 5(b3). In Fig. 5(b4), at energies $\epsilon>35$ the energy eigenstates tend to be also eigenstates of the operator $\hat{J}_{x}$, opening along the different values of $m_{x}^{2}$, which also explains that the expectation values of $\hat{J}_{z}$ in Fig. 5(b3) tend to zero.
A similar initial trend in $\langle r\rangle$ is observed in Fig. 5(c1), where it has large values at low energies before converging toward the Poisson limit. However, at higher energies, $\langle r\rangle$ begins to rise again, signaling the onset of spectral correlations. This behavior aligns with the increasing disorder observed in the corresponding Peres lattices [see Figs. 5(c2)–5(c4), column view], marking the beginning of a transition toward nonintegrability.
Figures 5(d1) and 5(e1) show that the average spectral ratio $\langle r\rangle$ fluctuates around the value expected for GOE statistics in certain energy regions. This behavior is consistent with the strong disorder observed in the corresponding Peres lattices at high energies [see Figs. 5(d2)–5(d4) and 5(e2)–5(e4), column view].
With the exception of the case $\gamma=0.01$, where the energy spectrum remains nearly harmonic and thus obscures the spectral statistics, we observe again a strong correspondence between the degree of disorder in the Peres lattices and the values of the average spectral ratio $\langle r\rangle$. As the coupling strength $\gamma$ increases, a clear trend emerges in which the spectral ratio transitions from values characteristic of Poisson statistics to those indicative of GOE statistics, reflecting a progressive loss of integrability and the emergence of quantum chaos.